2.) You are interested in calculating the sample error and size of the upcoming local elections. You asked 1000 people if they are planning on voting for candidate A. 40% said “Yes”. Calculate the margin of sample error at 95%. What sample size will you need to have a sample error of only 1% at the 95% confidence level?
Answer #1

(A)Solution :

Given that,

n = 1000

Point estimate = sample proportion = р = 0.40

1 - р = 1 - 0.40 = 0.60

At 95% confidence level

\alpha = 1 - 95%

\alpha =1 - 0.95 =0.05

\alpha/2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

Margin of error = E = Z\alpha / 2 * \sqrt ((р * (1 - р )) / n)

= 1.96 (\sqrt((0.40*0.60) /1000 )

E = 0.0304

(B)

Solution :

Given that,

\hat p = 0.40

1 - \hat p = 1 - 0.40= 0.60

margin of error = E = 1% = 0.01

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

Sample size = n = (Z\alpha/2 / E)2 * \hat p * (1 - \hat p )

= (1.96 / 0.01)2 * 0.40* 0.60

= 9220

Sample size = 9220